Fama’s Llamas has a weighted average cost of capital of 10.2 percent. The company’s cost of equity is 12 percent, and its pretax cost of debt is 7.1 percent. The tax rate is 24 percent. What is the company’s target debt-equity ratio? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., 32.1616.)
Answer #1

After-tax cost of debt=7.1*(1-tax rate)

=7.1*(1-0.24)=5.396%

Debt-equity ratio=debt/equity

Let debt be $x

Equity be $y

Total=$(x+y)

WACC=Respective costs*Respective weight

10.2=(x/(x+y)*5.396)+(y/(x+y)*12)

10.2*(x+y)=5.396x+12y

10.2x+10.2y=5.396x+12y

x=(12-10.2)y/(10.2-5.396)

=0.3747 y(Approx).

Hence debt-equity ratio=debt/equity

=0.3747(Approx).